# Generating Exceptional Behavior Tests with Reasoning Augmented Large Language Models

### v2

@@ VI-C RQ3: Selection of non-EBTs
In addition to randomly choosing one relevant non-EBT to add to the prompt for exLóng, we try prompting exLóng with no more than 5 different relevant non-EBTs (one prompt for each non-EBT) and reporting the best performance, dubbed exLóng-sample in tables IV and V. We see that that performance differences can be substantial.
TABLE VI: Ablations on different context of exLóng.
@@
exLóng 64.17 68.60 85.78 19.82 82.49 100.00 67.97 58.53
No stack trace 64.38 69.16 86.12 18.20 82.49 100.00 66.82 56.91
No stack trace, guard expression 63.61 68.08 85.64 18.43 80.41 100.00 60.60 49.31
No stack trace, guard expression, non-EBT 57.17 61.79 82.30 13.59 62.21 100.00 47.93 37.79
TABLE VII: Comparison between using different non-EBTs to prompt and using one non-EBT but sampling multiple times. Note that we add the target EBT’s name to the prompt and we only report results on examples that have more than one candidate non-EBT.
@@
exLóng 62.81 67.29 85.87 15.02 85.77 100.00 66.40 55.34
exLóng sample w. one non-EBT 63.83 69.13 86.86 14.62 87.75 100.00 73.12 62.06
exLóng sample w. different non-EBT 72.26 75.60 90.87 13.83 94.07 100.00 81.42 71.54
TABLE VIII: Results for exLóng-base 13B and exLóng 13B models; exLóng-base 13B does not have access to non-EBTs, stack trace and guard expression.
@@
exLóng-base 13B 57.89 63.40 82.52 13.59 65.90 100.00 53.46 43.09
exLóng 13B 64.44 68.77 85.93 19.82 80.65 99.43 69.59 62.21
To study how the diversity of non-EBTs used in the prompt affect the exLóng performance, our ablation study on the number of different relevant non-EBTs is shown in Table VII. We present results of 1) exLóng only sampling one EBT for each target throw statement; 2) exLóng with the same non-EBT as 1) but sample more than one EBT for each target throw statement; 3) exLóng sampling one time for at most five different non-EBTs (exLóng-sample). For a fair comparison, we keep the number of sampled EBTs for 2) and 3) same and we only report metrics on a subset of data on which we extracted more than one relevant non-EBT.
We find that increasing the number of sampled EBTs improves the performance when compared to only sampling one EBT, which can be attributed to more randomness when the sampling size is small. However, sampling with a diverse non-EBTs further boost the performance across most of the metrics under the same sample size. This motivates a new research question on how to choose better non-EBTs to the prompt for EBT generation, which we leave for future work.
VI-D RQ4: Transferring Findings to Different LLM Size
To explore if our design and contributions are transferable to a model of a different size, we show the results for a variant of exLóng that is built on CodeLlama-13B. We introduce one ablation model called exLóng-base where the stack trace, guard expression, and non-EBT are excluded from model’s input. Table VIII shows the performance of exLóng-13B-base and exLóng-13B, and we find that the benefits brought by our proposed context are consistent across different model sizes, i.e., 7B and 13B. This shows that combining program analysis technique with LLMs is promising and necessary to further boost the LLMs when generating EBTs regardless of the number of parameters in a model.
@@ VI-E RQ5: Machine-Oriented Use Case
TABLE IX: Throw Statements Coverage Rate for exLóng, Randoop and EvoSuite.
@@
Figure 3: Venn diagram that shows target throw statements coverage by exLóng, Randoop, and EvoSuite.
In Table IX, we present the throw statement coverage rate of exLóng-sample and two analysis-based test generation tools. EvoSuite could generate tests for all projects, while Randoop could not generate tests for three projects. We inspected the issues and found that: (1) Randoop crashed on OpenNMS/newts because this project kept throwing runtime exceptions (all related to com.codahale.metrics.ScheduledReporter); (2) Randoop crashed on pinterest/secor because this project requires the configuration of Kafka; (3) Randoop could not load a class and crashed on OpenHFT/Chronicle-Map.
We report results on the subset of 27 projects where all the tools can be run successfully (Subset Projects) and results on all 30 projects (All Projects). Among the given target throw statements, exLóng achieves higher throw statement coverage rate than analysis-based tools. Figure 3 illustrates the overlap and difference among the sets of target throw statements covered by exLóng, EvoSuite, and Randoop. All three tools cover different sets of throw statements. exLóng covers the most target throw statements that other two cannot. In this experiment, for exLóng, we sample only a single EBT per given prompt, and we expect that our results would be even better if with a larger sample size.
@@ VII Case Study
We submitted a subset of EBTs generated by exLóng to open-source repositories. We already submitted pull requests (PRs) to 9 repositories on GitHub, and 23 generated EBTs have been accepted by developers. In one instance, a developer responded and merged our PR only 30 minutes after we create the PR. This was encouraging, and future tool development should integrate exLóng into an IDE, such that exLóng continuously provide EBTs for code that a developer is editing. We plan to submit PRs to other projects in the coming weeks.
@@ VIII Limitations
Not every MUT has a destination test file. We leave the problem on finding or generating destination test file for any given method under test as future work.
LLMs. We built exLóng around CodeLlama [39], a recent open-source model. We believe that building on CodeLlama provides reproducibility guarantees that will help us and others to build on this work. Our contributions are the task, definition of a context for the task, tools for extracting the context, instruction fine-tuned model, and extensive evaluation. We expect that building exLóng on LLMs with similar architecture as CodeLlama would lead to similar results.
@@ IX Related Work
There has been significant work on test generation [36, 32, 12, 11, 55, 44]. We cover related work on (1) LLM-based test generation, (2) generating tests for exceptional behavior, and (3) other test generation techniques.
@@
Conditions are useful for guiding the generation of tests and finding bugs [4, 3, 40]. SymPrompt [40] introduced path constraint prompting to guide LLMs to generate high-coverage tests without additional training. Their path constraints were generated by parsing the abstract syntax tree of the target method while our guard expressions are derived by searching backwards from the throw statement to the top of target method that contains this throw statement.
@@
Generating tests for exceptional behavior. Exception handling [62, 61, 26, 63] is an important aspect of software development. There are several techniques [51, 14, 2, 1] to generate EBTs. However, they either generate EBTs from specifications or use random-based or search-based strategies to generate EBTs. Our work is the first to use LLMs to generate EBTs. Also, they generate tests for the whole program, while our exLóng allows users to specify which throw statements to cover.
@@
Other test generation techniques. Other techniques incorporate random-based [32, 38], search-based [12, 18] and constraint-based [11, 20, 13] strategies to automatically generate tests. Tests can be derived from multiple sources, including the code under test, error messages [17], and specifications [27] like comments [48, 46, 47].
@@ X Conclusion
We presented the first work on generating tests for exceptional behavior (EBTs) using large language models. We introduced exLóng that builds on top of CodeLlama to embed reasoning about traces that lead to throw statements, conditional expressions along those traces, and non-exceptional tests that cover similar traces. We evaluated exLóng in two use cases: developer-oriented use case (i.e., generate EBT for a given method under test and a target throw statement) and machine-oriented use case (i.e., automatically generate tests for all throw statements available in a repository). Our results show that exLóng outperforms existing test generation models and analysis-based test generation tools. We contributed a number of tests generated by exLóng to open-source projects, and 23 EBTs are already accepted. We believe that exLóng targets an important task, has good performance, and helps developers increase code quality assurance by automatically providing high quality EBTs.
@@ Acknowledgments
We thank Nader Al Awar, Yu Liu, Jayanth Srinivasa, Aditya Thimmaiah, Zijian Yi and Zhiqiang Zang for their comments and feedback. We acknowledge the Texas Advanced Computing Center (TACC) at The University of Texas at Austin and Digital Research Alliance of Canada (the Alliance) for providing HPC resources that have contributed to the research results reported within this paper. This work is partially supported by the US National Science Foundation under Grant Nos. CCF-2107291, CCF-2217696 and CCF-2313027. This work was in part supported by Cisco Research. Any opinions, findings and conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of Cisco Research.

### v3

@@ VI-C RQ3: Selection of non-EBTs
In addition to randomly choosing one relevant non-EBT to add to the prompt for exLong, we try running the inference of exLong for no more than 5 times, each time with a different relevant non-EBT, and reporting the best performance, dubbed exLong-sample in tables IV and V. We see that that performance differences can be substantial.
To study how the diversity of non-EBTs used in the prompt affect the exLong performance, our ablation study on the number of different relevant non-EBTs is shown in Table VII. We present results for 1) exLong generating only one EBT for each target throw statement; 2) exLong-sample but using the same non-EBT as 1) in multiple inference runs; 3) exLong-sample using different non-EBTs in multiple inference runs. For a fair comparison, we keep the number of generated EBTs for 2) and 3) same and we report the best-of-$k$ metrics on a subset of 253 examples on which we extracted more than one relevant non-EBT.
We find that increasing the number of sampled EBTs improves the performance when compared to only generating one EBT, which can be attributed to the randomness. Moreover, sampling with a diverse set of non-EBTs further boost the performance across most of the metrics under the same sample size. This motivates a new research question on how to choose better non-EBTs to the prompt for EBT generation, which we leave for future work.
VI-D RQ4: exLong Built with GPT-4o
TABLE VI: Ablations on different context of exLong.
@@ VI-D RQ4: exLong Built with GPT-4o
exLong 63.13 67.49 85.32 19.05 82.10 100.00 67.36 59.45
No stack trace 62.61 67.36 85.43 17.74 81.41 100.00 67.05 58.14
No stack trace, guard expression 62.53 67.42 84.85 17.74 81.41 100.00 61.98 52.00
No stack trace, guard expression, non-EBT 54.84 60.62 80.70 12.06 61.52 100.00 46.70 38.25
TABLE VII: Comparison between using different non-EBTs to sample and using the same non-EBT but sampling multiple times. Note that we add the target EBT’s name to the prompt and we only report results on examples that have more than one candidate non-EBT.
@@
exLong 61.44 66.21 85.34 13.44 83.40 100.00 66.14 56.79
exLong-sample w/ same non-EBT 63.64 68.28 86.70 13.18 87.35 100.00 72.46 61.53
exLong-sample w/ different non-EBT 71.66 75.18 90.62 17.79 95.65 100.00 83.00 71.94
TABLE VIII: Comparison between GPT-4o-few-shot and exLong–GPT-4o.
@@
GPT-4o-few-shot 60.07 65.56 84.00 16.82 81.87 99.73 71.52 55.53
exLong–GPT-4o 60.48 66.77 84.77 17.74 82.49 100.00 75.35 64.75
Our technique, which assists LLMs in generating EBTs can be transferred to other LLMs with significant benefits. In addition to results with open-sourced LLMs in prior sections, we further present the results of exLong built on GPT-4o [35], currently one of the most powerful proprietary LLMs. Table VIII compares the performance of exLong–GPT-4o with baseline GPT-4o under the developer-oriented use case. The inputs to GPT-4o includes the MUT, the target exception type, the method containing the target throw statement, one relevant non-EBT, and the destination test file. We instruct exLong–GPT-4o with the prompt containing MUT, stack trace, guard expression, relevant non-EBTs and the destination test file. One example pair of input and ground-truth EBT is provided for both models. Results show that exLong–GPT-4o outperforms GPT-4o across both similarity metrics and functional-correctness metrics.
@@ VI-E RQ5: Machine-Oriented Use Case
TABLE IX: Throw statements coverage rate for exLong, Randoop and EvoSuite.
@@
Figure 4: Venn diagram that shows target throw statements coverage by exLong, Randoop, and EvoSuite.
In Table IX, we present the throw statement coverage rate of exLong-sample and two analysis-based test generation tools. EvoSuite could generate tests for all projects, while Randoop could not generate tests for three projects. We inspected the issues and found that: (1) Randoop crashed on $\mathtt{OpenNMS/newts}$ because this project kept throwing runtime exceptions (all related to $\mathtt{com.codahale.metrics.ScheduledReporter}$); (2) Randoop crashed on $\mathtt{pinterest/secor}$ because this project requires the configuration of Kafka; (3) Randoop could not load a class and crashed on $\mathtt{OpenHFT/Chronicle\text{-}Map}$.
We report results on the subset of 27 projects where all the tools can be run successfully (Subset Projects) and results on all 30 projects (All Projects). Among the given target throw statements, exLong achieves higher throw statement coverage rate than analysis-based tools. Figure 4 illustrates the overlap and difference among the sets of target throw statements covered by exLong, EvoSuite, and Randoop. All three tools cover different sets of throw statements. exLong covers the most target throw statements that other two cannot.
@@ VII Case Study
We performed a case study where we submitted the EBTs generated by exLong to the open-source projects (where the data was extracted from) to collect developers’ feedback. Among the evaluation set for machine-oriented use case, exLong generated 187 EBTs across 30 projects that are runnable and cover the correct throw statements. We selected a subset of 9 projects that are actively maintained, i.e., they had at least one commit, accepted pull request (PR) or responded issue within the past six months (at the time of the paper submission). We found that the generated EBTs of 2 projects were the same as those added by developers on later commits (commits after the ones we used during evaluation), thus refrained from submitting PRs to them. In total, we submitted 7 PRs which include 35 EBTs (one PR per project). Among them, 4 PRs (23 EBTs) have been accepted, and 3 PRs (12 EBTs) are still pending. No PR was rejected. In one instance, a developer responded and merged our PR only 30 minutes after we create the PR. This was encouraging, and future tool development should integrate exLong into an IDE, such that exLong continuously provide EBTs for code that a developer is editing.
@@ VIII Limitations
Destination test file. Not every MUT has a destination test file. We leave the problem on finding or generating destination test file for any given method under test as future work.
LLMs. We built exLong around CodeLlama [43], a recent open-source model. We believe that building on CodeLlama provides reproducibility guarantees that will help us and others to build on this work. Our contributions are the task, definition of a context for the task, tools for extracting the context, instruction fine-tuned model, and extensive evaluation. We expect that building exLong on other open-source LLMs would lead to similar results.
@@ IX Related Work
There has been significant work on test generation [40, 36, 11, 10, 59, 48] and code generation [67, 68, 38, 30, 27, 23, 33]. We cover related work on (1) LLM-based test generation, (2) generating tests for exceptional behavior, and (3) other test generation techniques.
@@
Conditions are useful for guiding the generation of tests and finding bugs [4, 3, 44]. SymPrompt [44] introduced path constraint prompting to guide LLMs to generate high-coverage tests without additional training. They collect constraints from each possible execution path in the target method and prompt the LLM to generate tests that cover those paths. We extract guard expressions by analyzing multiple methods along the stack trace starting from the throw statement to the target method for generating EBTs.
@@
Generating tests for exceptional behavior. Exception handling [66, 65, 28, 69] is an important aspect of software development. There are several techniques [55, 13, 2, 1] to generate EBTs. However, they either generate EBTs from specifications or use random-based or search-based strategies to generate EBTs. Our work is the first to use LLMs to generate EBTs. Also, prior work generates tests for the whole program, while our exLong allows users to specify which throw statements to cover.
@@
Other test generation techniques. Other techniques incorporate random-based [36, 42], search-based [11, 17, 25, 26] and constraint-based [10, 19, 12] strategies to automatically generate tests. Tests can be derived from multiple sources, including the code under test, error messages [16], and specifications [29] like comments [52, 50, 51].
@@ X Conclusion
We presented the first work on generating tests for exceptional behavior (EBTs) using large language models. We introduced exLong that builds on top of CodeLlama to embed reasoning about traces that lead to throw statements, conditional expressions along those traces, and non-exceptional tests that cover similar traces. We evaluated exLong in two use cases: developer-oriented use case (i.e., generate EBT for a given method under test and a target throw statement) and machine-oriented use case (i.e., automatically generate tests for all throw statements available in a repository). Our results show that exLong outperforms existing test generation models and analysis-based test generation tools. We contributed a number of tests generated by exLong to open-source projects, and 23 EBTs are already accepted. We believe that exLong targets an important task, has good performance, and helps developers increase code quality assurance by automatically providing high quality EBTs.
@@ Acknowledgments
We thank Nader Al Awar, Jayanth Srinivasa, Aditya Thimmaiah, Zijian Yi, Samuel Yuan, Zhiqiang Zang, Linghan Zhong, and the anonymous reviewers for their comments and feedback. This work is partially supported by the US National Science Foundation under Grant Nos. CCF-2107291, CCF-2217696, CCF-2313027, CCF-2403036; as well as AST-2421782 and Simons Foundation MPS-AI-00010515 (NSF-Simons AI Institute for Cosmic Origins – CosmicAI). This work was in part supported by Cisco Research. Any opinions, findings and conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of Cisco Research.